OpenStudy (agentjamesbond007):

Derivative of x^3cos3x^2

OpenStudy (anonymous):

3x^2cos(3x^2)-6x^4sin(3x^2)x

OpenStudy (agentjamesbond007):

@jamalahmed2068, can you explain how you got that answer

OpenStudy (anonymous):

do you know product rule??? (uv)'=u'v+uv' just use this..

OpenStudy (agentjamesbond007):

Yes. Do you need to use the chain rule?

OpenStudy (anonymous):

yes you have to use chain rule for the derivative of cos(x^2)

OpenStudy (agentjamesbond007):

How come? Is it necessary for any Trig. function drivative?

OpenStudy (anonymous):

y=cos(x^2) let u=x^2 du=2xdx y=cos(u) dy=-sin(u)du so dy=-2xsin(x^2)dx

OpenStudy (anonymous):

not always!! chain rule dy/du=(dy/dx) (dx/du) got it!!

OpenStudy (anonymous):

chain rule is used when functions are composed....(do you know the the composition of functions)

OpenStudy (agentjamesbond007):

Is it the f(g(x)) type

OpenStudy (anonymous):

f(x)=cos(x) g(x)=x^2 fog=cos(x^2)

OpenStudy (anonymous):

let f and g are functions of x y=f(g(x)) let u=g y=f(u) dy/dx=d(f(u))/du du/dx =f'(u)du/dx =f'(g(x)) g'(x) got it nowww